An object falls from a high building and hits the ground in 9.0 seconds. Ignoring air resistance, what is the distance that it fell?


1
Expert's answer
2022-05-05T13:57:00-0400
"H=\frac{gt^2}{2}=\frac{9.8*9.0^2}{2}=397\:\rm m"